# VesselSAM: Leveraging SAM for Aortic Vessel Segmentation with LoRA and Atrous Attention

### v1

@@ III-B Preliminary: SAM architecture
The SAM[6] is a prompt-based segmentation framework composed of three main components: the Image Encoder, Prompt Encoder, and Mask Decoder. The Image Encoder is based on a ViT, which processes input images using 16 × 16 pixel patches through transformer blocks to capture image features, resulting in image embedding. The Prompt Encoder handles various prompts, including points, bounding boxes and masks, converting them into feature vectors that guide the segmentation. The Mask Decoder is a two-layer transformer-based decoder that fuses image embedding and prompt features using cross-attention. It incorporates a Multi-Layer Perceptron (MLP) for feature refinement and dimensionality alignment, and utilizes convolutional layers for upsampling to produce high-resolution masks.
@@ III-C VesselSAM
The VesselSAM architecture builds on the foundation of SAM, with several key modifications aimed at improving vascular structure segmentation in medical images. As illustrated in Fig. 1, VesselSAM incorporates the Atrous Attention module and LoRA layers, designed to capture multi-scale features and reduce the number of trainable parameters while maintaining segmentation accuracy.
@@ III-D LoRA and AtrousLoRA
LoRA [17] has emerged as a PEFT method, enabling task-specific adaptations of pre-trained models while significantly reducing computational and memory overhead. LoRA introduces low-rank trainable matrices to approximate weight updates, effectively bypassing the need to fine-tune the entire model(Fig. 2(a)). Instead, it adds two small matrices, $W_{e}$ and $W_{d}$, while keeping the original weights $W_{0}$ frozen during training. Given a pre-trained weight matrix $W_{0}\in\mathbb{R}^{C_{\text{out}}\times C_{\text{in}}}$, LoRA modifies the forward pass of the model as
$h=W_{0}x+W_{d}W_{e}x,$ (1)
where $W_{0}$ is the frozen pre-trained weight matrix, $W_{e}\in\mathbb{R}^{r\times C_{\text{in}}}$ and $W_{d}\in\mathbb{R}^{C_{\text{out}}\times r}$ are the low-rank encoder and decoder matrices, and $r$ is the rank of the decomposition, with $r\ll\min(C_{\text{in}},C_{\text{out}})$. Here, $x\in\mathbb{R}^{B\times C_{\text{in}}}$ represents the input, where $B$ is the batch size.
@@
Mathematically, with AtrousLoRA, Equation 1 changes to:
$h=W_{0}x+W_{d}\cdot\text{Atrous}(W_{e}x),$ (2)
@@
where $W_{0}\in\mathbb{R}^{C_{\text{out}}\times C_{\text{in}}}$ is the frozen pre-trained weight matrix, $W_{e}\in\mathbb{R}^{r\times C_{\text{in}}}$ and $W_{d}\in\mathbb{R}^{C_{\text{out}}\times r}$ are the low-rank encoder and decoder matrices, and $x\in\mathbb{R}^{B\times C_{\text{in}}\times H\times W}$ is the input feature map. In this case, $B$ represents the batch size, $C_{\text{in}}$ and $C_{\text{out}}$ are the input and output channels, and $H$ and $W$ represent the height and width of the feature maps. The Atrous module applies atrous convolutions with predefined dilation rates to $W_{e}x$, effectively capturing multi-scale contextual features. Atrous LoRA employs fixed dilation rates, which simplifies implementation while maintaining adaptability for various vision-related tasks.
AtrousLoRA retains the efficiency of LoRA while extending its applicability to tasks requiring spatial and contextual understanding, such as semantic segmentation and medical image analysis. The predefined dilation rates ensure a balance between computational efficiency and multi-scale feature extraction.
@@
III-E Atrous Attention Module
We propose a novel attention mechanism for vision transformers, called Atrous Attention Module (Fig. 2(c)), which performs a fusion between regional and sparse attention. This approach allows us to capture both global context and local detail with efficient computational complexity, while preserving the hierarchical information present in medical images. Inspired by atrous convolution [24], which increases the receptive field by skipping rows and columns in the input feature map without adding extra parameters, Atrous Attention enables VesselSAM to focus on relevant anatomical structures across multiple scales. The process is shown in Algorithm 1.
The data flow within the Atrous Attention Module starts by passing the input feature map $X\in\mathbb{R}^{B\times C\times H\times W}$ through the Atrous Spatial Pyramid Pooling (ASPP) [21], which applies dilated convolutions at different rates $d_{i}$ to capture features at various scales. Each atrous convolution produces an output feature map $Y_{i}=f(X;W_{i},d_{i})$ where $W_{i}$ are the convolution weights and $d_{i}$ is the dilation rate. Additionally, global average pooling is performed on $X$ to obtain $Z=f_{\text{1x1}}(\text{GAP}(X))$ The outputs from ASPP, including the atrous convolutions $Y_{i}$ and the global pooling result $Z$, are concatenated into a Concatenated Feature Map $Y=[Y_{1},Y_{2},\dots,Y_{n},Z]$. This concatenated feature map then goes through a 1x1 Convolution, reducing it to the desired number of output channels, followed by Batch Normalization (BN) and ReLU activation, generating the ASPP Output $Y_{\text{ASPP}}=\text{ReLU}(\text{BN}(f_{\text{1x1}}(Y)))$.
@@ III-E Atrous Attention Module
III-F Atrous Spatial Pyramid Pooling
Atrous Spatial Pyramid Pooling (ASPP) is another important component in the VesselSAM Model. It plays a crucial role in enhancing the model’s ability to segment vascular structures by capturing multi-scale contextual information from medical images. VesselSAM, which integrates advanced segmentation techniques, utilizes ASPP to improve the segmentation of blood vessels by applying dilated convolutions with varying dilation rates. This allows the model to capture both fine details and broader contextual information without losing resolution. ASPP increases the receptive field by using dilated convolutions at multiple rates, enabling the model to understand both local features of vessels and their spatial relationships within the image, which is critical for accurate vascular segmentation.
Mathematically, in VesselSAM, ASPP begins by applying dilated convolutions with different dilation rates $d_{i}\in\{d_{1},d_{2},\dots,d_{n}\}$, where each rate captures features at different scales. For each dilation rate $d_{i}$, the dilated convolution operation is applied to the input feature map $X$ as follows:
$Y_{i}=f_{\text{dil}}(X;W_{i},d_{i})=X*_{d_{i}}W_{i},$ (3)
where $*_{d_{i}}$ denotes the dilated convolution, and $W_{i}$ represents the filter with dilation rate $d_{i}$. In addition to the dilated convolutions, a global average pooling operation is applied to capture the global context of the input image, which is mathematically defined as:
$Z=\frac{1}{H\times W}\sum_{h=1}^{H}\sum_{w=1}^{W}X_{b,c,h,w},$ (4)
where $Z$ represents the global pooling result, effectively reducing the spatial dimensions to $1\times 1$ while maintaining the channel information. These outputs are then concatenated into a single feature map:
$Y_{\text{concat}}=[Y_{1},Y_{2},\dots,Y_{n},Z],$ (5)
The concatenated feature map $Y_{\text{concat}}$ contains multi-scale information, helping VesselSAM capture both local vessel features and broader contextual relationships, which is essential for accurately segmenting complex vascular structures. To reduce the dimensionality of this concatenated feature map, a $1\times 1$ convolution is applied:
$Y_{\text{ASPP}}=f_{\text{1x1}}(Y_{\text{concat}})=W_{\text{1x1}}\cdot Y_{\text{concat}}+b_{\text{1x1}},$ (6)
where $W_{\text{1x1}}$ and $b_{\text{1x1}}$ are the weight and bias of the $1\times 1$ convolution. Finally, a non-linear activation function, such as ReLU, is applied to introduce non-linearity into the model:
$Y_{\text{ASPP}}=\text{ReLU}(Y_{\text{ASPP}})$ (7)
In VesselSAM, ASPP is integral to capturing multi-scale contextual features, allowing the model to handle objects, such as blood vessels, at different scales effectively. By using dilated convolutions with various rates, ASPP enables VesselSAM to segment vascular structures accurately while maintaining computational efficiency. This multi-scale feature extraction is essential for addressing the variability in vessel sizes and complex vascular networks commonly encountered in medical imaging.
@@ III-G Prompt Encoder And Mask Decoder
In VesselSAM, the Prompt Encoder remains frozen, ensuring the stability of the pre-trained parameters while allowing for efficient processing of user prompts. In our case, the prompts are provided in the form of bounding boxes, which are represented by their top-left and bottom-right corner points. Each corner point is mapped into a 256-dimensional embedding, which serves as the input to the segmentation process. By freezing the prompt encoder, VesselSAM enables real-time interaction, as the image embedding can be precomputed, allowing the user to provide bounding-box input dynamically without the need for retraining.
On the other hand, the Mask Decoder in VesselSAM is fully trainable and plays a crucial role in producing the segmentation output. The decoder architecture includes two transformer layers, which are responsible for fusing the image embedding with the prompt embeddings through cross-attention. This fusion allows the bounding box information to guide the segmentation task effectively. Following this, the decoder employs two transposed convolution layers to upsample the combined embedding to a resolution of 256 × 256, ensuring a high level of detail is retained in the final segmentation mask. The output is then passed through a sigmoid activation function, followed by bi-linear interpolation to match the resolution of the original input image, thereby producing the final high-resolution mask.
@@
Note: Bold indicates the best results and underline denotes the second best results.
@@
Note: Bold indicates the best results and underline denotes the second best results.
@@ IV-A Datasets
In our experiments, we utilized two key datasets to evaluate the effectiveness of the proposed VesselSAM model in complex medical segmentation tasks. The Aortic Vessel Tree (AVT) Segmentation dataset [27] comprises 56 contrast-enhanced CT angiography (CTA) scans collected from three sources: the KiTS Grand Challenge, the Rider Lung CT dataset, and Dongyang Hospital. Among these, 38 cases were designated for training, while the remaining 18 were used for testing. All slices were resampled to a spatial resolution of 1 mm × 1 mm, with Hounsfield Unit (HU) values normalized to [0, 1]. Additionally, the TBAD dataset [28], comprising 100 CTA images from Guangdong Provincial People’s Hospital, was utilized for segmenting True Lumen (TL), False Lumen (FL), and False Lumen Thrombus (FLT) in Type-B Aortic Dissection (TBAD) cases. To conform to Segment Anything Model (SAM) requirements, both the AVT and TBAD datasets were converted from 3D CTA volumes into 2D slices. Each 3D scan was converted into NumPy arrays, with all slices resampled to a uniform resolution of 1 mm × 1 mm. Voxel intensity values were normalized using standard CT window settings [400, 40]. Ground truth masks were refined by removing labels of irrelevant structures and small objects, using thresholds of 1000 voxels for 3D volumes and 100 pixels for individual 2D slices. Only non-zero slices were retained, and intensity normalization was applied. The processed 2D slices were then resized to 1024 × 1024 pixels and converted into three-channel images by duplicating the grayscale slice across three channels (1024 × 1024 × 3), ensuring compatibility with SAM’s input format.
IV-B Loss Function and Evaluation Metrics
We have used a combined loss function comprising an unweighted sum between cross-entropy loss and Dice loss, which has been widely adopted for its robustness in medical image segmentation tasks[10]. Let $P$ represent the predicted segmentation output, and $T$ denote the corresponding ground truth. For each voxel $j$, $p_{j}$ and $t_{j}$ correspond to the predicted and ground truth values, respectively. The total number of voxels in the image is denoted by $M$. The binary cross-entropy loss is defined as:
$L_{CE}=-\frac{1}{M}\sum_{j=1}^{M}[t_{j}\log p_{j}+(1-t_{j})\log(1-p_{j})],$ (8)
The final loss $L$ is computed as the sum of the cross-entropy loss and the Dice loss:
$L=L_{CE}+L_{D}$ (10)
To evaluate the performance of the segmentation model, we employed two widely used metrics: Dice Similarity Coefficient (DSC) and Hausdorff Distance (HD). The DSC measures the spatial overlap between the predicted segmentation $P$ and the ground truth $T$, and is defined as:
$DSC(P,T)=\frac{2|P\cap T|}{|P|+|T|}$ (11)
where $\partial P$ and $\partial T$ represent the boundaries of the predicted and ground truth regions, respectively, and $d(x,y)$ is the Euclidean distance between points $x$ and $y$. A lower HD value indicates better boundary alignment between the predicted and ground truth segmentations.
These evaluation metrics provide complementary insights into the model’s performance, with DSC focusing on region-based overlap and HD emphasizing boundary accuracy. Together, they offer a comprehensive assessment of the segmentation quality.
Figure 3: Qualitative visual results on the AVT (Dongyang, Rider, Kits) and TBAD Datasets under Bounding Box Prompts. The first column shows the input images, followed by the ground truth (GT) in the second column. The subsequent columns show the results of various segmentation models: VesselSAM, SAM, MedSAM, SAMAdp, SAMedAdp, and SAM-MedIA. Each model’s output is overlaid with color-coded regions for true lumen (pink), false lumen (green), and the GT boundary line (yellow). The blue box represents the bounding box-prompt used for segmentation. The images have been zoomed in to enhance visibility.

### v2

@@ III-B Preliminary: SAM architecture
The SAM[6] is a prompt-based segmentation framework composed of three main components: the Image Encoder, Prompt Encoder, and Mask Decoder. The Image Encoder is based on a ViT, which processes input images using 16 × 16 pixel patches through transformer blocks to capture image features, resulting in image embedding. The Prompt Encoder handles various prompts, including points, bounding boxes and masks, converting them into feature vectors that guide the segmentation. These prompt embeddings enable SAM to focus on specific regions of interest within an image, improving segmentation accuracy and adaptability. The Mask Decoder is a two-layer transformer-based module that fuses image embedding and prompt features using cross-attention mechanisms. To refine feature representations and ensure precise mask generation, the decoder incorporates a Multi-Layer Perceptron (MLP) for feature refinement and dimensionality alignment. Additionally, convolutional layers are utilized for upsampling, allowing the model to produce high-resolution segmentation masks.
@@ III-C VesselSAM
The VesselSAM architecture builds on the foundation of SAM framework, incorporating key modifications to improve aortic vessel segmentation. As illustrated in Fig. 1, VesselSAM integrates the Atrous Attention module and LoRA layers, designed to capture multi-scale features and reduce the number of trainable parameters while maintaining segmentation accuracy.
@@ III-D LoRA and AtrousLoRA
LoRA [17] has emerged as a PEFT method, enabling task-specific adaptations of pre-trained models while significantly reducing computational and memory overhead. LoRA introduces low-rank trainable matrices to approximate weight updates, effectively bypassing the need to fine-tune the entire model (Fig. 2 (a)). Specifically, it adds two small matrices, $W_{b}$ and $W_{a}$, while keeping the original weights $W_{O}$ frozen during training. Given a pre-trained weight matrix $W_{O}\in\mathbb{R}^{C_{\text{out}}\times C_{\text{in}}}$, LoRA modifies the forward pass of the model as
$y=W_{O}x+W_{b}W_{a}x$ (1)
where $W_{O}$ is the frozen pre-trained weight matrix, $W_{b}\in\mathbb{R}^{r\times C_{\text{in}}}$ and $W_{a}\in\mathbb{R}^{C_{\text{out}}\times r}$ are the low-rank encoder and decoder matrices, and $r$ is the rank of the decomposition, with $r\ll\min(C_{\text{in}},C_{\text{out}})$. Here, $x\in\mathbb{R}^{B\times C_{\text{in}}}$ represents the input, where $B$ is the batch size.
@@
Mathematically, with AtrousLoRA Eq. 1 changes to:
$y=W_{O}x+W_{b}\cdot\text{AtrousAttention}(W_{a}x)$ (2)
@@
where $W_{O}\in\mathbb{R}^{C_{\text{out}}\times C_{\text{in}}}$ is the frozen pre-trained weight matrix, $W_{a}\in\mathbb{R}^{r\times C_{\text{in}}}$ and $W_{b}\in\mathbb{R}^{C_{\text{out}}\times r}$ are the low-rank encoder and decoder matrices, and $x\in\mathbb{R}^{B\times C_{\text{in}}\times H\times W}$ is the input feature map. In this case, $B$ represents the batch size, $C_{\text{in}}$ and $C_{\text{out}}$ are the input and output channels, and $H$ and $W$ represent the height and width of the feature maps. The AtrousAttention module applies atrous convolutions with predefined dilation rates to $W_{a}x$, effectively capturing multi-scale contextual features. The AtrousAttention can be formulated as:
$\text{AtrousAttention}(W_{a}x)=Y_{\text{ASPP}}\odot A_{\text{sigmoid}}$ (3)
where $Y_{\text{ASPP}}$ is the output of the Atrous Spatial Pyramid Pooling (ASPP) module, and $A_{\text{sigmoid}}$ is the attention map generated by the attention mechanism. The element-wise product $Y_{\text{ASPP}}\odot A_{\text{sigmoid}}$ combines the multi-scale features from the ASPP module with the attention map, resulting in an attention-weighted feature map that enhances important regions and suppresses less relevant ones. This mechanism enables AtrousLoRA to focus on the most salient features while maintaining contextual information across multiple scales.
III-E Atrous Spatial Pyramid Pooling
Atrous Spatial Pyramid Pooling (ASPP), originally proposed by [21], capable of capturing multi-scale contextual information. In our work, ASPP is integrated into the Atrous Attention Module to enhance the segmentation of vascular structures in medical images. By leveraging dilated convolutions with varying dilation rates, ASPP enables the model to capture both fine details and broader contextual information without sacrificing resolution. This capability is particularly important for accurately segmenting blood vessels, as it allows the model to understand both local features and their spatial relationships within the image.
Mathematically, ASPP operates by applying dilated convolutions with multiple dilation rates $d_{i}\in\{d_{1},d_{2},\dots,d_{n}\}$, where each rate $d_{i}$ extracts features at a specific scale. For a given input feature map $X\in\mathbb{R}^{B\times C\times H\times W}$, the dilated convolution operation for each rate $d_{i}$ is defined as.
$Y_{i}=f_{\text{dil}}(X;W_{i},d_{i})=X*_{d_{i}}W_{i}$ (4)
where $*_{d_{i}}$ denotes the dilated convolution operation, and $W_{i}$ represents the convolutional filter with dilation rate $d_{i}$.
In addition to the multi-scale dilated convolutions, ASPP incorporates a global average pooling (GAP) operation to capture the global context of the input feature map. Mathematically the GAP operation is defined as:
$Z=\frac{1}{H\times W}\sum_{h=1}^{H}\sum_{w=1}^{W}X_{b,c,h,w}$ (5)
where $Z\in\mathbb{R}^{B\times C\times 1\times 1}$ represents the globally pooled feature map, summarizing the spatial information into a single vector per channel.
The outputs of the dilated convolutions $Y_{i}$ and the global average pooling $Z$ are then concatenated into a single feature map as expressed in Eq. 6.
$Y_{\text{concat}}=[Y_{1},Y_{2},\dots,Y_{n},Z]$ (6)
where $Y_{\text{concat}}\in\mathbb{R}^{B\times C^{\prime}\times H\times W}$ combines multi-scale features, enabling the model to capture both local and global contextual information. To reduce the dimensionality of the concatenated feature map, a $1\times 1$ convolution is applied.
$Y_{\text{ASPP}}=f_{\text{1x1}}(Y_{\text{concat}})=W_{\text{1x1}}\cdot Y_{\text{concat}}+b_{\text{1x1}}$ (7)
where $W_{\text{1x1}}$ and $b_{\text{1x1}}$ are the weight and bias of the $1\times 1$ convolution, respectively. Finally, a non-linear activation function ReLU [29] is applied.
$Y_{\text{ASPP}}=\text{ReLU}(Y_{\text{ASPP}})$ (8)
III-F Atrous Attention Module
The Atrous Attention Module (Fig. 2(c)) is introduced as a novel attention mechanism for Vision Transformers, designed to fuse regional and sparse attention effectively. This approach allows us to capture both global context and local detail with efficient computational complexity, while preserving the hierarchical information present in medical images. Inspired by atrous convolution [24], which expands the receptive field by skipping rows and columns in the input feature map without increasing the number of parameters. Atrous Attention enables VesselSAM to focus on relevant anatomical structures across multiple scales. The process is shown in Algorithm 1.
The data flow within the Atrous Attention Module starts by passing the input feature map $X\in\mathbb{R}^{B\times C\times H\times W}$ through the ASPP, which applies dilated convolutions at different rates $d_{i}$ to capture features at various scales. Each atrous convolution produces an output feature map $Y_{i}=f(X;W_{i},d_{i})$ where $W_{i}$ are the convolution weights and $d_{i}$ is the dilation rate.
@@ III-F Atrous Attention Module
[…]
@@ III-G Prompt Encoder And Mask Decoder
In VesselSAM, the Prompt Encoder remains frozen, ensuring the stability of the pre-trained parameters while allowing for efficient processing of user prompts. In our case, the prompts are provided in the form of bounding boxes, which are represented by their top-left and bottom-right corner points. Each corner point is mapped into a 256-dimensional embedding, which serves as the input to the segmentation process. By freezing the prompt encoder, VesselSAM enables real-time interaction. The image embedding can be precomputed, allowing users to dynamically provide bounding-box input without the need for retraining.
On the other hand, the Mask Decoder in VesselSAM is fully trainable and plays a crucial role in producing the segmentation output. The decoder architecture includes two transformer layers, which are responsible for fusing the image embedding with the prompt embeddings through cross-attention. This fusion allows the bounding box information to guide the segmentation task effectively. The mask decoder employs two transposed convolution layers to upsample the combined embedding to a resolution of 256 × 256 while ensuring a high level of detail is retained in the final segmentation mask. The output is then passed through a sigmoid activation function followed by bi-linear interpolation to match the resolution of the original input image thereby producing the final high-resolution mask.
III-H Loss Function and Evaluation Metrics
We have used a combined loss function comprising an unweighted sum between cross-entropy loss and Dice loss, which has been widely adopted for its robustness in medical image segmentation tasks[10]. As detailed in Eq.(9), Eq.(10), and Eq.(11), where $P$ represents the predicted segmentation output and $T$ denote the corresponding ground truth. For each voxel $j$, $p_{j}$ and $t_{j}$ correspond to the predicted and ground truth values, respectively. The total number of voxels in the image is denoted by $M$. The binary cross-entropy loss is defined as:
$L_{CE}=-\frac{1}{M}\sum_{j=1}^{M}[t_{j}\log p_{j}+(1-t_{j})\log(1-p_{j})],$ (9)
The final loss $L$ is computed as the sum of the Dice loss and the cross-entropy loss :
$L=L_{Dice}+L_{CE}$ (11)
To evaluate the performance of the segmentation model, we employed two metrics: Dice Similarity Coefficient (DSC) and Hausdorff Distance (HD). The DSC measures the spatial overlap between the predicted segmentation $P$ and the ground truth $T$, and is defined as:
$DSC(P,T)=\frac{2|P\cap T|}{|P|+|T|}$ (12)
where $\partial P$ and $\partial T$ represent the boundary points of the predicted and ground truth regions respectively and $d(x,y)$ is the Euclidean distance between points $x$ and $y$. A lower HD value indicates better boundary alignment between the predicted and ground truth segmentations.
@@ III-H Loss Function and Evaluation Metrics
Note: Bold indicates the best results and underline denotes the second best results. ”Vanilla” refers to versions using pre-trained weights, while ”FT” indicates fine-tuned versions on AVT datasets.
@@
Note: Bold indicates the best results and underline denotes the second best results. ”Vanilla” refers to versions using pre-trained weights, while ”FT” indicates fine-tuned versions on TBAD dataset.
@@ IV-A Datasets
In our experiments, we utilized two key datasets to evaluate the effectiveness of the proposed VesselSAM model in complex medical segmentation tasks. The Aortic Vessel Tree (AVT) Segmentation dataset [33] comprises 56 contrast-enhanced CT angiography (CTA) scans collected from three sources: the KiTS Grand Challenge, the Rider Lung CT dataset, and Dongyang Hospital. Among these, 38 cases were designated for training, while the remaining 18 were used for testing. All slices were resampled to a spatial resolution of 1 mm × 1 mm, with Hounsfield Unit (HU) values normalized to [0, 1]. Additionally, the TBAD dataset [34], comprising 100 CTA images from Guangdong Provincial People’s Hospital, was utilized for segmenting True Lumen (TL), False Lumen (FL), and False Lumen Thrombus (FLT) in Type-B Aortic Dissection (TBAD) cases. To conform to the SAM requirements, both the AVT and TBAD datasets were converted from 3D CTA volumes into 2D slices. Each 3D scan was converted into NumPy arrays, with all slices resampled to a uniform resolution of 1 mm × 1 mm. Voxel intensity values were normalized using standard CT window settings [400, 40]. Ground truth masks were refined by removing labels of irrelevant structures and small objects, using thresholds of 1000 voxels for 3D volumes and 100 pixels for individual 2D slices. Only non-zero slices were retained, and intensity normalization was applied. The processed 2D slices were then resized to 1024 × 1024 pixels and converted into three-channel images by duplicating the grayscale slice across three channels (1024 × 1024 × 3), ensuring compatibility with SAM’s input format.
IV-B Implementation Details
All experiments were conducted using the VesselSAM model implemented with the PyTorch deep learning library. VesselSAM is based on the SAM architecture, employing a ViT-Base image encoder initialized with pre-trained weights. During training, the parameters of the image encoder and prompt encoder remained frozen, while fine-tuning was applied exclusively to the mask decoder and the integrated AtrousLoRA modules. The AtrousLoRA module comprises ASPP and attention mechanisms. Specifically, the ASPP utilized dilated convolutions with dilation rates of 1, 6, 12, and 18, enabling the capture of multi-scale contextual information critical for accurate segmentation. Additionally, LoRA layers with a low-rank dimension of 4 were integrated to achieve an optimal trade-off between model accuracy and computational efficiency by reducing the number of trainable parameters to approximately 7% of the total parameters.
The model was optimized using the AdamW optimizer with an initial learning rate of 1e-4 and a weight decay of 0.01. Training proceeded for a total of 100 epochs with a batch size of 8. To further improve computational efficiency, mixed-precision training was employed. Data augmentation techniques included random perturbations to bounding box coordinates to improve the model’s generalizability. All experiments were performed on a single NVIDIA H100 GPU with 80GB VRAM. The model was evaluated on two challenging benchmark datasets: the Aortic Vessel Tree (AVT) and the Type-B Aortic Dissection (TBAD). Performance metrics included the Dice Similarity Coefficient (DSC) and the Hausdorff Distance (HD).